Look at this function: g(x)= x^3+x^2-x+1 What is the value of g(e)= ?


Look at this function: g(x)= x^3+x^2-x+1 What is the value of g(x+e)+?

Answer by algebrapro18(206)

well for the first one just plug in a e for x

g(x)= x^3+x^2-x+1
g(e)= e^3+e^2-e+1

for the secound one just plug in x+e for x

g(x)= x^3+x^2-x+1
g(x+e)= (x+e)^3+(x+e)^2-(x+e)+1